# RealRAG: Retrieval-augmented Realistic Image Generation via Self-reflective Contrastive Learning

### v2

@@ 3 Methodology
Problem Setting: Given a textual prompt $T$ drawn from a text space $\mathcal{T}$, the text-to-image generation task aims to produce a corresponding image $I\in\mathcal{I}$ that accurately reflects the semantics of $T$. Formally, we model this task as learning a conditional distribution $p(I\mid T)$, which describes how likely an image $I$ is given the text $T$. In practice, we often parameterize this distribution with a neural generator $G_{\theta}$ (with parameters $\theta$), yielding:
@@ 3.1 Self-reflective Contrastive Learning
where $N$ is the training batch size, $\mathcal{D}_{neg}^{nor}$ is the similarity between the text prompt and the in-batch negative, $\mathcal{D}_{neg}^{ref}$ is the similarity between the text prompt and the reflective negative.
@@ 3.2 Real-object-based Retrieval-augmented Generation
where $Retriever$ is the trained reflective retriever, and $n$ in the scale of real-object-base database.
@@ 4.2.2 Generative Result
Qualitative Results. We show the qualitative results in Fig. 4. The visual results show the ability of our RealRAG to overcome hallucinations and significantly improve the realism and quality of fine-grained realistic image generation. Specifically, for the autoregressive model, In the prompt [”Audi TT Hatchback 2011”], the RealRAG model produces a highly realistic rendering of the vehicle, capturing intricate details like the side mirrors and headlights, which are closer to the real object compared to the original AR model with significant hallucination in the car’s side. For the U-Net-based diffusion model, In prompt [”Silverbush”], RealRAG ensures the correct flower structure and vibrant color tones, whereas the original model struggles with maintaining fine-grained details like petal shape. Lastly, compared with the original generator, RealRAG delivers a compelling representation of the breed’s signature coat patterns and body posture from the [”Bluetick”] prompt. The results demonstrate that RealRAG generates high realism and correct objects and show the flexible application of the SoTA generators.

### v3

@@ 3 Methodology
Problem Setting: Given a textual prompt $T$ drawn from a text space $\mathcal{T}$, the text-to-image generation task aims to produce a corresponding image $I\in\mathcal{I}$ that accurately reflects the semantics of $T$. Formally, we model this task as learning a conditional distribution $p(I\mid T)$, which describes how likely an image II is given the text $T$. In practice, we often parameterize this distribution with a neural generator $G_{\theta}$ (with parameters $\theta$), yielding:
@@ 3.1 Self-reflective Contrastive Learning
where NN is the training batch size, $\mathcal{D}_{neg}^{nor}$ is the similarity between the text prompt and the in-batch negative, $\mathcal{D}_{neg}^{ref}$ is the similarity between the text prompt and the reflective negative.
@@ 3.2 Real-object-based Retrieval-augmented Generation
where RetrieverRetriever is the trained reflective retriever, and $n$ in the scale of real-object-base database.
@@ 4.2.2 Generative Result
Qualitative Results. We show the qualitative results in Fig. 4. The visual results show the ability of our RealRAG to overcome hallucinations and significantly improve the realism and quality of fine-grained realistic image generation. Specifically, for the autoregressive model, In the prompt [”Audi $T$ Hatchback 2011”], the RealRAG model produces a highly realistic rendering of the vehicle, capturing intricate details like the side mirrors and headlights, which are closer to the real object compared to the original AR model with significant hallucination in the car’s side. For the U-Net-based diffusion model, In prompt [”Silverbush”], RealRAG ensures the correct flower structure and vibrant color tones, whereas the original model struggles with maintaining fine-grained details like petal shape. Lastly, compared with the original generator, RealRAG delivers a compelling representation of the breed’s signature coat patterns and body posture from the [”Bluetick”] prompt. The results demonstrate that RealRAG generates high realism and correct objects and show the flexible application of the SoTA generators.